Accelerator Update for December 10, 2008 - December 12, 2008
Wednesday Dec 10
The day shift began with Operations monitoring the Tevatron (TeV) store 6631 (162.64E30) and the Recycler stash (100.7E10), with the Antiproton Source stacking (34.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the MiniBooNE experiment taking beam. NuMI was down due to a horn problem.
At 1:38 PM, the P1 transfer line beam permit tripped off when a power supply’s current (I:V701) drifted low. Everything reset.
Pbar experts reported at 2:15 PM that they had set a New Stacking Record: 27.35 mA/hour.
Operations reported at 3:57 PM that the Recycler had stashed antiprotons eight times during the day shift.
Operations established beam to the Meson M01 area absorber at 4:42 PM.
Operations established beam to MTest at 5:32 PM.
Operations disabled the MTest beam at 6:40 PM.
Operations terminated store 6631 at 6:43 PM.
Operations put the TeV into shot setup at 6:55 PM.
At 7:58 PM, Operations reported a MI RF station (MIRF5) that was tripping off. Later, after talking with an expert, operators bypassed the station.
Operations reported at 8:15 PM that NuMI and safety experts had completed a special two-hour run to test and inspect the newly installed horn.
Operations established store 6632 at 8:24 PM with an initial luminosity of 324.7E30.
A TeV expert began a Tevatron Electron Lens (TEL) study at 8:30 PM.
Operators began preparing NuMI for access at 9:25 PM.
The TEL study ended at 11:07 PM.
Operations reported at midnight that the Recycler had stashed eight times during the evening shift.
Thursday Dec 11
The midnight shift began with Operations monitoring store 6632 (172.23E30) and the stash (109.5E10), with Pbar stacking (30.1E10), and with MiniBooNE taking beam.
At 5:07 AM, an Accumulator RF station (ARF1-2) tripped off for the first of two times in 2.5 hours. It reset.
Operations reported at 8:04 AM that the Recycler had stashed eight times during the midnight shift.
Operations established beam to MTest at 8:47 AM.
Operations terminated store 6632 at 11:45 AM.
Operations put the TeV into shot setup at noon.
Operations established store 6633 at 1:37 PM with an initial luminosity of 328.68E30.
At 10:53 PM, Booster quad B:Q802 tripped on phase unbalance, an expert investigated.
Friday Dec 12
The midnight shift began with Operations monitoring store 6633 (102.7E30) and the stash (227.8E10), with Pbar in standby (30.3E10), MiniBooNE waiting for beam.
B:Q802 was back on at 12:14 AM
Beam to Pbar and MiniBooNE at 12:19 AM.
Operations terminated store 6633 at 6:32 AM.
Operations put the TeV into shot setup at 7:01 AM.
The Plans for Friday and the Weekend
The plans for today are to stack, stash, and store, NuMI horn repairs are complete and NuMI should be starting up this afternoon.
Comments and Suggestions
|last modified 01/05/2009 email Fermilab|
|Security, Privacy, Legal|